OpenStudy (anonymous):

Help me find the derivative of this equation.....show steps: y = 6(1-(t/12))² y' = ?

OpenStudy (anonymous):

\[6(1-(t/12))^{2}=6((12-t)/12)^{2}\]\[(6/144)(12-t)^{2}=(1/24)(144-24t+t ^{2})\]\[=(144-24t+t ^{2})/24\]. Hence y'= 0-24/24 +2t/24 =-1+t/12